### v1

# REDUCIO! Generating 1024$\times$1024 Video within 16 Seconds using Extremely Compressed Motion Latents

Abstract
Commercial video generation models have exhibited realistic, high-fidelity results but are still restricted to limited access. One crucial obstacle for large-scale applications is the expensive training and inference cost. In this paper, we argue that videos contain much more redundant information than images, thus can be encoded by very few motion latents based on a content image. Towards this goal, we design an image-conditioned VAE to encode a video to an extremely compressed motion latent space. This magic Reducio charm enables 64$\times$ reduction of latents compared to a common 2D VAE, without sacrificing the quality. Training diffusion models on such a compact representation easily allows for generating 1K resolution videos. We then adopt a two-stage video generation paradigm, which performs text-to-image and text-image-to-video sequentially. Extensive experiments show that our Reducio-DiT achieves strong performance in evaluation, though trained with limited GPU resources. More importantly, our method significantly boost the efficiency of video LDMs both in training and inference. We train Reducio-DiT in around 3.2K training hours in total and generate a 16-frame 1024$\times$1024 video clip within 15.5 seconds on a single A100 GPU. Code released at https://github.com/microsoft/Reducio-VAE.
[…]
[width=autoplay=True,keepaspectratio]8figs/vsample_1/000015
Impressionist style, a yellow rubber duck floating on the waves at sunset.
[…]
[width=autoplay=True,keepaspectratio]8figs/vsample_2/000015
An aerial shot of a lighthouse standing tall on a rocky cliff, its beacon cutting through the early dawn, waves crashing against the rocks below.
A wizard waves his wand.
Figure 1: Examples of generated high-resolution videos. Click the top row images to play the video clips. Best viewed with Adobe Reader.
1 Introduction
The straightforward but surprisingly dangerous charm, Reducio, cause certain things to shrink.
Miranda Goshawk   “Harry Potter” [37]
Recent advances in latent diffusion models (LDMs) [36] for video generation have presented inspiring results [55, 4, 1, 3, 47, 58, 2, 5, 32], showing great potential in various applications. Commercial models like Sora [27], Runway Gen-3 [38], Movie Gen [31], and Kling [17] can already generate photorealistic and high-resolution video clips. However, training and deploying such models are computationally demanding—thousands of GPUs and millions of GPU hours are required for training, and the inference for a one-second clip costs several minutes. Such a high cost is becoming a significant obstacle for research and large-scale real-world applications.
Extensive effort has been made to relieve the computational burden by means of efficient computation modules in the backbone [44, 54], optimizing diffusion training strategies [11, 51], and adopting few-step sampling methods [41, 57]. However, in this paper, we argue that the very nature of the problem is long neglected. Concretely, most video LDMs stick to the paradigm of text-to-image diffusion, e.g., Stable Diffusion (SD) [36], and inherently employ the latent space of its pre-trained 2D variational autoencoder (VAE) [16, 34]. This scheme compresses the input videos by 8×8$\times$8 × in each spatial dimension, which is suitable for images while being much surplus for video.
As videos naturally carry much more redundancy than images, we believe they can be projected into a much more compressed latent space with a VAE based on 3D convolutions. Surprisingly, for videos, a 16×16$\times$16 × spatial down-sampling factor does not sacrifice much of the reconstruction performance. While the resulting distribution of latent space shifts from those of the pre-trained image LDMs, we can still wield their well-learned spatial prior by factorizing text-to-video generation with a condition frame [10, 48, 53, 61]. In particular, we can obtain a content frame with off-the-shelf text-to-image LDMs and then generate samples with text and image as joint priors into video LDMs.
More importantly, as the image prior delivers enriched spatial information of video contents, factorized video generation offers the potential for further compression. We argue that video can be encoded into very few latents representing motion variables, i.e., motion latents, plus a content image. Hereby, we introduce the magical Reducio charm by building a VAE comprised of a 3D encoder that aggressively compresses input videos into a 4096×4096$\times$4096 × down-sampled compact space, and a 3D decoder fused with feature pyramids of the middle frame as the content condition. Notably, the resulting Reducio-VAE surpasses the common 2D VAE by 5db in PSNR while using a $\mathbf{64\times}$ smaller latent space.
Subsequently, we establish the LDM with a diffusion transformer (DiT) [28, 6]. Besides using T5 features [33] as the text condition, we employ an image semantic encoder as well as a context encoder to provide additional image conditions, which inform the model of the spatial content. Thanks to the extremely compressed video latent, our diffusion model, namely Reducio-DiT, thus enjoys both the fast training (inference) speed and high generation quality. Specifically, Reducio-DiT achieves 318.5 FVD score on UCF-101, surpassing a bunch of previous work [4, 1, 49, 62, 48] while being much faster. It can also be easily upscaled to a larger resolution, e.g. $1024^{2}$, with affordable cost. Our experiments show that Reducio-DiT can generate a 16-frame video clip of $1024^{2}$ resolution within 15.5 seconds on a single A100 GPU, and hence achieves a $\mathbf{16.6\times}$ speedup over Lavie [49].
In summary, our contributions are as follows: (1) We encode video into very few motion latents plus a content image, where the designed Reducio-VAE can compress video into a $\mathbf{64\times}$ smaller latent space compared to common 2D VAE. (2) Based on Reducio-VAE, a diffusion model called Reducio-DiT is devised by incorporating additional image conditions. (3) Experiments show that our Reducio-DiT can significantly speed up the generation process while producing high-quality videos.
2 Related Work
Figure 2: Architecture and performance comparisons between vanilla 2D VAE in (a) SDXL, (b) 3D VAE and (c) the proposed Reducio-VAE. Reducio-VAE enables 64$\times$ reduction of latent size while achieving much higher reconstruction performance.
Video diffusion models have become the latest craze in the field of text-to-video generation. Due to the huge success of text-to-image LDMs [12, 24, 40, 23, 36], attempts [4, 3, 47] have been made to leverage the powerful UNet-based stable diffusion models pre-trained in 2D space by inserting temporal convolution and attention layers. Recently, the strong performance of transformers has continued in image generation domain [8, 6, 18] by diffusion transformers (DiT) [28]. Consequently, recent works have focused on extending DiT-based diffusion models from 2D spatial attention to spatiotemporal attention.
While the mainstream exploration [31] follows the paradigm of image diffusion models by using text as the only condition, an emerging stream of works [10, 48, 61] decompose text-to-video diffusion into two steps, i.e., text-to-image and text-image-to-video generation. This factorizing encourages models to concentrate on motion modeling instead of challenging spatiotemporal joint modeling. In this work, we take advantage of factorized text-to-video generation and build our framework upon DiT. Thanks to the delicate structure of Reducio-VAE, our model manages to preserve the fine details of image priors in generated videos, and hence can be used for more high-demanding real-world applications, e.g., advertisement generation.
Efficient diffusion models have gained increasing attention due to the tremendous computational resources and time required for training and inferring samples, especially for high-resolution images and videos. On one hand, researchers have made attempts to build LDMs with more efficient architectures [9, 44, 52, 64]. For example, DIM [44] uses Mamba as its backbone to ease the computational burden compared to vanilla transformers. Another line of work leverages efficient diffusion techniques to perform fewer sampling steps during inference [21, 41, 57] or to accelerate convergence during training [11, 51]. For instance, Stable Diffusion 3 [8] and Flux [18] employ rectified flow [19] for improved training and inference efficiency.
Our work focuses on designing a more compact latent space [29], which straightforwardly and effectively reduces the overall training and inference costs of LDM. PVDM [59] and CMD [60] propose projecting 3D video samples into three separate yet compact 2D latent spaces. However, this approach requires training separate diffusion models for each space, which inevitably leads to discrepancies during inference. Our Reducio-VAE takes after LaMD [14] to adopt a factorized, compact motion latent space with the help of a content frame for reconstruction. In contrast, we emphasize aggressive compression of the spatial dimension and suggest that a moderate temporal down-sampling factor provides the best trade-off between performance and efficiency. Concurrent research in the image domain also indicates that using a high spatial down-sampling factor on redundant visual content [43, 52], such as high-resolution images, significantly improves efficiency. We propose that it is crucial to recognize the importance of spatial redundancy in the video latent space.
3 Reducio
Figure 3: The overview of Reducio-DiT training framework. Compared to DiT, Reducio-DiT requires an additional image-condition module to inject the semantics and content of the keyframe via cross-attention.
3.1 Reducio Video Autoencoder
Variational autoencoder (VAE) [16] projects input images or videos into a compressed latent space obeying a certain distribution and maps the latent sampled from the obtained distribution into RGB space with a decoder. Latent diffusion models [36] take advantage of the down-sampled latent space of VAE to support generation with improved efficiency. In this paper, we follow the common practice and build our video autoencoder upon VAE.
We denote the temporal down-sampling factor of a VAE as $f_{t}$ and the spatial one as $f_{s}$. Given an input video $\mathcal{V}$ with the shape of $3\times T\times H\times W$, the compressed latent $z_{\mathcal{V}}$ as the input into diffusion models fall into the space of $|z|\times T/f_{t}\times H/f_{s}\times W/f_{s}$, where $z$ denotes the latent channel. Thanks to the superior performance of open-source Stable Diffusion (SD) [36] models, it has been a common practice to leverage SD-VAE to process images and videos into compressed latent. Typically, as shown in (a) of Fig. 2, SD-VAE is built upon 2D convolution with $f_{s}=8$ and $f_{t}=1$.
Reducio charm on videos. As adjacent video frames share large similarities in pixel space, video signals are inherently more redundant than images. Based on this idea, we adopt a more aggressive down-sampling factor for video autoencoder. Specifically, we inflate the 2D convolutions into 3D convolutions and increase the number of autoencoder blocks, increasing $f_{t}$ and $f_{s}$ to 4 and 16, as illustrated in (b) of Fig. 2. While the overall down-sampling factor increases by 16×16$\times$16 ×, the compressed latent space suffices to achieve comparable pixel-space reconstruction performance, making a strong case for the potential of extremely shrinking video latent dimension.
As for image-to-video generation, we argue that video latent can be further compressed as the content image contains rich visual details and substantial overlapping information. In such a scenario, video latents can be further simplified to represent the motion information. We can find similar spirits in video codec [35], where carefully designed modules use reference frames as inputs to perform high-fidelity reconstruction of a given compressed code. Specifically, we propose to cast the Reducio charm on video latent space, as shown in (c) of Fig. 2. Generally, we choose the middle frame $V_{T/2}$ as the content guidance, compress input videos by an extreme down-sampling factor with a 3D encoder, and reconstruct from the latent space with the aid of content frame features pyramids. Specifically, we infuse content features of sizes varying in $H/8\times W/8$ and $H/4\times W/4$ via cross-attention within the decoder, adopt an extreme compression ratio, i.e., $f_{s}=32$ and $f_{t}=4$. While Reducio-VAE projects videos into 64×64$\times$64 × more compressed latent space than one of SDXL-VAE [30], it significantly outperforms the latter in reconstruction performance.
3.2 Reducio Diffusion Transformer
Thanks to the magical compression capability of Reducio-VAE, we project input videos into compressed latent space with a down-sampling factor of 4096, and hence significantly speed up the training and inference of diffusion models. In specific, we adopt the DiT-XL [28] model and follow most of the design choices adopted by PixArt-$\alpha$ [6], i.e., AdaLN-single modules, cross attention layers with Flan-T5-XXL [33] text conditions. The training target follows the reverse diffusion process [12], which can be depicted as
$\bm{p}_{\theta}(\bm{z}_{t-1}|\bm{z}_{t})=\mathcal{N}(\bm{z}_{t-1};\bm{\mu}_{\theta}(\bm{z}_{t},t),\bm{\sigma}_{t}^{2}\mathbf{I}),$ (1)
where $\bm{z}$ is latent, $t$ is time step, $\bm{\mu}_{\theta}(\bm{z}_{t},t)$ and $\bm{\sigma}_{t}^{2}$ indicate the mean and variance of the sample at the current time step. By re-parameterizing $\bm{\mu}_{\theta}$ as a noise prediction network $\varepsilon_{\theta}$, the model can be trained using simple mean-squared error between the predicted noise $\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)$ and the ground truth sampled Gaussian noise $\bm{\varepsilon}$ formulated as
$\mathcal{L}_{simple}=\mathbb{E}_{\bm{z}_{0},\bm{\varepsilon},t}(\left\|\bm{\varepsilon}-\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)\right\|^{2}).$ (2)
To adapt the image diffusion model to video, we consider two options: (1) Directly transforming the 2D attention to full 3D attention, without adding additional parameters. (2) Adding temporal layers and transforming the model to perform 2D spatial attention plus 1D temporal attention. We adopt option (1) by default. Studies on the evaluation of these two options are shown in Sec. 4.3.2.
In addition, to tailor for image-conditioned video generation, we introduce Reducio-DiT (illustrated in Fig. 3) with additional image-condition modules, as described below.
Content frame modules consist of a semantic encoder built upon pretrained OpenCLIP ViT-H [7] and a content encoder initialized with the SD2.1-VAE [36]. The former encoder projects content frames into a high-level semantic space while the latter one mainly focuses on extracting the spatial information. We concatenate the obtained features with text tokens output by T5 to form the image-text-joint condition, and then cross-attend them with the noisy video latent. Concretely, the detailed operations can be described as:
$\bm{e}_{s}=\mathtt{OpenCLIP}(V_{T/2}),$ (3)
$\bm{e}_{c}=\mathtt{SD\mbox{-}VAE}(V_{T/2}),$
$\bm{e}_{img}=\mathtt{MLP}(\bm{e}_{s})\oplus\mathtt{MLP}(\bm{e}_{c}),$
$\bm{e}_{p}=\mathtt{MLP}(\mathtt{T5}(prompt)),$
$\bm{e}=[\bm{e}_{img},\bm{e}_{p}],$
$\bm{z}_{i}=\mathtt{CAttn}(\bm{z}_{i},\bm{e}),$
where $V_{T/2}$ is the condition image, $\oplus$ refers to element-wise addition, []delimited-[][$\cdot$][ ⋅ ] refers to concatenation, and $\mathtt{CAttn}(\cdot)$ refers to cross-attention.
Table 1: Quantitative comparison with open-source state-of-the-art VAE for generation.
Model Downsample $|\bm{z}|$ Pexels UCF-101
[…]
SD2.1-VAE [36] $1\times 8\times 8$ 4 29.23 0.82 0.09 25.96 21.00
SDXL-VAE [30] $1\times 8\times 8$ 16 30.54 0.85 0.08 19.87 23.68
OmniTokenizer [46] $4\times 8\times 8$ 8 27.11 0.89 0.07 23.88 30.52
OpenSora-1.2 [26] $4\times 8\times 8$ 16 30.72 0.85 0.11 60.88 67.52
Cosmos-VAE [25] $8\times 8\times 8$ 16 30.84 0.74 0.12 29.44 22.06
$8\times 16\times 16$ 16 28.14 0.65 0.18 77.87 119.37
Reducio-VAE $4\times 32\times 32$ 16 35.88 0.94 0.05 17.88 65.17
Figure 4: Comparison between the $1^{st}$ and $16^{th}$ frames of videos reconstructed by SDXL-VAE (top) and Reducio-VAE (bottom).
Scaling up to high-resolution videos consumes tremendous computational resources for vanilla video diffusion models. Nevertheless, Reducio-DiT relieves this limit to a great extent. To support high-resolution (e.g., $1024^{2}$) video generation, we employ a progressive training strategy: 1) in the first stage of training, the model learns to align the video latent space with text-image prior by taking in a vast number of $256^{2}$ videos as input; 2) in the second stage, we fine-tune the model on videos with higher resolution, i.e., $512\times 512$. In consequence, the content encoder augments 4×4$\times$4 × more tokens with the noisy latent; 3) in the third stage, we conduct further fine-tuning on videos with spatial resolution around $1024$ with multi-aspect augmentation.
To better accommodate the model to videos with varying aspect ratios, we inject the aspect ratio and size of input videos as embeddings into the DiT model. Then we add the dynamic size embedding with timestep embedding before being fed into the MLP. More importantly, for models in the first and the second stages, the content encoder cross-attends $H/16\times W/16$ tokens with the noisy latent. As computational costs grow rapidly when resolution increases, we introduce a strategy to ensure the efficiency of Reducio-DiT on high-resolution videos. Specifically, we aims to shrink the content embedding by 2×2$\times$2 × in the spatial dimension in each cross-attention layer. Inspired by Deepstack [22], we divide input tokens into 4 groups of grids and iteratively concatenate each group of $H/32\times W/32$ tokens with the other condition tokens.
4 Experiments
4.1 Training and Evaluation Details
Our Reducio-VAE is trained on a 400K video dataset collected from Pexels***https://www.pexels.com/. Pexels contains a large number of free and high-quality stock videos, where each video is accompanied with a short text description. We train Reducio-VAE from scratch on $256\times 256$ videos with 16 FPS. Note that we split videos with higher resolutions into tiles in the spatial dimension. To perform the VAE comparison, PSNR, SSIM [50], LPIPS [63] and reconstruction FVD (rFVD) [45] are employed as the evaluation metrics.
Reducio-DiT is trained on the above Pexels dataset and an internal dataset with 5M videos containing high-resolution text-video pairs. As described in Sec. 3.2, we adopt a multi-stage training strategy to train the model from low resolution to high resolution. We first train Reducio-DiT-256 using a batch size of 512 on 4 Nvidia A100 80G GPUs for around 900 A100 hours and initialize model weights with PixArt-$\alpha$-256 [6]. Then we fine-tune on $512^{2}$ videos shortly for 300 A100 hours to obtain Reducio-DiT-512. In the third stage, we randomly sample video batches from 40 aspect ratio buckets, which is the same as the setting in PixArt-$\alpha$. We leverage 8 AMD MI300 GPUs to support a training batch size of 768 and fine-tune Reducio-DiT-1024 for 1000 GPU hours.
Table 2: Quantitative comparison with state-of-the-art video LDMs.
[…]
UCF-101 MSR-VTT $256\times 256$ $1024\times 1024$ A100 Training
Make-A-Video [39] - - 367.23 - - - -
CogVideo [13] 7.8B 67.01 701.59 1294 0.0023 OOM $>$368K
Show-1 [62] 1.7B 78.93 394.46 538.0 0.0016 OOM $>$56K
Lavie [49] 1B 77.08 526.30 - 0.1838 0.0039 -
MicroCinema [48] 2B - 342.86 377.4 0.0234 OOM $>$25K
Lumiere [2] $>$2B - 332.49 - - - -
Reducio-DiT 1.2B 81.39 318.50 291.9 0.9824 0.0650 3.2K
A child excitedly swings on a rusty swing set, laughter filling the air.
Figure 5: Comparison between frames generated given an identical frame and prompt, by (a) DynamicCrafter [53], (b) SVD-XT [3] and (c) Reducio-DiT, respectively. We resize the output frames from $1344\times 768$ to $1024\times 576$ to match with the former two baselines.
We adopt DPM-Solver++ [20] as the sampling algorithm for efficient inference and set the sampling step to 20. For evaluation in Tab. 2, we calculate the speed of generating 16-frame video clips on a single A100 80G GPU and denote out-of-memory as OOM. For Reducio-DiT, we use PixArt-$\alpha$-256 and PixArt-$\alpha$-1024 to generate the content frame for $256^{2}$ and $1024^{2}$ videos, respectively. We report the FVD and IS scores of Reducio-DiT-512 under the zero-shot setting on UCF-101 [42] and MSR-VTT [56]. Since our method is a two-stage video generation approach, which requires a condition image prior, we follow MicroCinema [48] to use SDXL [30] to generate the condition image for all the evaluations, including the FVD on UCF-101 and MSR-VTT. Additionally, we evaluate Reducio-DiT-512 on the recent video generation benchmark VBench [15].
4.2 Main Results
Reducio-VAE preserves fine details of video inputs. Fig. 4 displays the first and the last frame of $256^{2}$ videos reconstructed by SDXL-VAE and Reducio-VAE, with the yellow bounding box highlighting the region with obvious differences in details. While SDXL-VAE causes corruptions and blurs, Reducio-VAE generally maintains the subtle texture within the middle frame. Tab. 1 also reflects the advantage of Reducio-VAE in quantitative reconstruction metrics, e.g., PSNR, SSIM, and LPIPS on a randomly sampled validation subset of 1K videos with the duration of 1s on Pexels, and test on both UCF-101 and Pexels for FVD.
As shown in the table, Reducio-VAE strikes obvious advantages on overall reconstruction performance. Specifically, our model significantly outperforms state-of-the-art 2D VAE, e.g., SD2.1-VAE and SDXL-VAE, by more than 5 db in PSNR. Moreover, Reducio-VAE also demonstrates better performance compared with VAEs designed for video in recent literature, e.g., OmniTokenizer and OpenSora-1.2. Our VAE also outperforms the concurrent work, Cosmos-VAE, which also adopts a further compressed video latent space, by 0.2 in SSIM and 5 db in PSNR with an 8×8$\times$8 × higher down-sampling factor.
We observe Reducio-VAE has worse rFVD than others on UCF-101. By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence result in a lower recontsruction quality. In contrast, Reducio-VAE places first on rFVD on pexels, which consists videos with much improved viusal quality and much higher resolution.
Reducio-DiT achieves a promising balance between efficiency and performance. As we harness an extremely compressed latent space, the training cost of our diffusion models is similar to the one of the text-to-image diffusion model. With a training cost of merely 3.2K A100 hours, our model achieves FVD scores of 318.50 on UCF-101 and 291.0 on MSR-VTT and the VBench score of 81.39, beating a range of previous stat-of-the-art video LDMs, as recorded in Tab. 2. Fig. 5 displays the comparison between Reducio-DiT and the other two stat-of-the-art image-to-video models. While DynamicCrafter and SVD-XT both fail to generate stable frames consistent with the given content frame, Reducio-DiT yields high visual quality with facial details in the content frame uncorrupted.
Moreover, Reducio-DiT significantly boost the throughput of video LDMs. The compressed latent space facilitates Reducio-DiT to output $256^{2}$ videos of 16 frames in almost 1 second, which realize a $\mathbf{5.3\times}$ speed up compared with Lavie. Moreover, 16-frame $1024^{2}$ video generation costs Reducio-DiT only for 15.4 seconds on average, leading to a striking speedup of $\mathbf{16.6\times}$ over Lavie. We argue the core philosophy of Reducio lies in the extremely compressed latent space, which acts as the key role in accelerating both training and inference speed of video LDMs.
4.3 Ablation Studies
4.3.1 Design of Reducio-VAE
Spatial redundancy is the key for compression. As Tab. 3 displays, when adopting an identical overall down-sampling factor $\mathcal{F}=f_{t}\cdot f_{s}^{2}$, Reducio-VAE achieves worse performance with larger compression factor in the temporal dimension. The result suggests that when a reference frame is provided, videos contain heavier redundancy in the spatial dimension than in the temporal dimension. Similarly, when $f_{s}=32$ is kept unchanged, increasing $f_{t}$ from 2 to 4 leads to a drop in PSNR by 0.91 db. To balance the trade-off between more compact latent space and better reconstruction performance, we set $f_{t}=4$ and $f_{s}=32$ in default.
Table 3: Ablation on the down-sampling factor of Reducio-VAE.
[…]
      16       16       34.14       0.93
      2       32       36.79       0.96
      4       32       35.88       0.94
      1       64       35.51       0.93
Scaling up latent channel is helpful for improving the reconstruction performance, as results shown in Tab. 4. When $\bm{z}$ grows from 4 to 16, the PSNR of reconstructed videos increases by 2.71 db. However, increasing the latent channel has a ceiling effect, as setting $\bm{z}$ to 32 hardly achieves further gains in both PSNR and SSIM. Therefore, we opt to use latent space with $|\bm{z}|=16$ for Reducio as the default setting. A visual qualitative comparison is presented in Fig. 6, where we compare the results when $|\bm{z}|=16$, $8$, and $4$.
Table 4: Ablation on latent channel dimensions of Reducio-VAE.
[…]
        4         33.17         0.91
        8         34.15         0.93
        16         35.88         0.94
        32         35.24         0.94
Figure 6: Comparison between the $8^{\text{th}}$ to $11^{\text{th}}$ frames of videos reconstructed by Reducio-VAE with different latent space channel dimensions, i.e., (a) $|\bm{z}|=16$, (b) $|\bm{z}|=8$, (c) $|\bm{z}|=4$.
Fusion with cross-attention achieves the most competitive performance in Tab. 5. As for add-based fusion, we feed the content frame features into a convolution and replicate the output on temporal dimension. Then we add video features with the content feature, together with a temporal embedding. Concerning linear-based fusion, we view the content condition as a single-frame video feature and concatenate it with the video feature by the temporal dimension. The concatenated features are then refined by a 3D convolution layer. As observed in Tab. 5, the attention-based Reducio-VAE outperforms other baselines by 1.71 in PSNR and 0.01 in SSIM, while demanding higher computational costs. By default, we display results using the attention-based fusion. Note that all Reducio-VAE displayed in Tab. 5 share the same 3D encoder and hence an identical latent space, where users are free to choose any of the decoders with different resource constraints.
Table 5: Ablation on the fusion type in Reducio-VAE when infusing the content image.
[…]
w/o. 27.91 0.80 435M
linear 30.86 0.87 439M
add 34.17 0.93 450M
attention 35.88 0.94 437M
4.3.2 Design of Reducio-DiT
For ablation studies on DiT, we conduct training for 300 A100 hours for each experiment on Pexels and report FVD and IS scores on UCF-101.
Using patch size of one yields the best performance. As shown in Tab. 6, keeping a fixed spatial down-sampling factor of 32, we compare the performance of using a 2$\times$ higher temporal down-sampling factor in latent space and a temporal patch size of 1 ($f_{t}=4$, $p_{t}=1$), with temporal patch size of 2 when patchify the input latents for obtaining patch embedding ($f_{t}=2$, $p_{t}=2$). We observe that compressing the temporal dimension with VAE outperforms. Similarly, we ablate using VAE with 2$\times$ lower spatial down-sampling factor yet setting the patch size of spatial dimension to 2 ($f_{s}=16$, $p_{s}=2$). The experimental results identify that compressing the video with Reducio-VAE to an extreme extent and using a small patch size, i.e., 1, in both the temporal and spatial dimensions achieves better performance.
Table 6: Ablation on the temporal patch size $p_{t}$ and the spatial patch size $p_{s}$ in Reducio-DiT.
[…]
2 32 2 1 351.0 32.7
4 16 1 2 534.1 30.4
4 32 1 1 337.6 34.1
Using joint spatiotemporal attention (3D attention) outweighs using factorized spatial and temporal attention (i.e., 2D + 1D attention) in generation quality. Interestingly, we observe that factorized attention leads to a faster convergence of training loss. However, with the same training step, as shown in Tab. 7, factorized attention lags behind its counterpart with joint 3D attention for 45 in FVD. We suppose the possible reason is that 2D + 1D scheme demands adding additional temporal layers and performs factorized self-attention on a small set of tokens e ach, making it hard to model smooth open-set motion with the light computation. In contrast, 3D attention directly exploits the original parameters and collaborates all spatiotemporal tokens.
A stunning sight as a steam train leaves the bridge, traveling over the arch-covered
viaduct. The landscape is dotted with lush greenery and rocky mountains.
Figure 7: Comparison on conditioning type in Reducio-DiT. (a) using semant.+content features. (b) using semantic features only.
Table 7: Ablation on the attention type in Reducio-DiT.
[…]
       2d + 1d        382.2        32.4
       3d        337.6        34.1
Fusion with both semantic and content information contributes to the best quality. As shown in Tab. 8, using semantic features, i.e., OpenClip-based features alone may contribute to distorted visual detail, leading to a lower FVD. on the other hand, using content-based features helps to make video smooth and stable, achieving a higher FVD yet less versatile motion. A comparison example is illustrated in Fig. 7. The collaboration of content and semantic features helps Reducio strike the best balance between content consistency and motion richness.
Table 8: Ablation on the conditioning type in Reducio-DiT.
[…]
     semantic      472.1      34.5
     content      363.7      31.1
     semantic + content      337.6      34.1
5 Conclusion
Video generation has shown promising results with many potential applications, but is still trapped by the unaffordable computation cost. In this paper, we explored how to effectively cut the overhead by reducing the latent size. Particularly, we found the video can be encoded into extremely compressed latents with the help of a content image prior, where the latent codes only need to represent the motion variables. With this observation, we designed the Reducio-VAE to compress videos to 4096×4096$\times$4096 × smaller latents. Using this powerful Reducio-VAE, we trained the Reducio-DiT that enables fast high-resolution video generation. Our method is also compatible with other accelerating techniques, e.g. rectified flow, allowing for further speedup, where we leave it for future exploration.
Limitations and future work. While Reducio-DiT shows a strong capacity of being consistent with the center frame, the generated videos are still relatively short (16 frames of 16 fps, i.e., 1 second) for real-world applications. Given the length of the video, the magnitude of the motion is also somehow limited in generated videos. We believe in the potential of adapting our works for longer video generation and will explore it as the future work.

### v2

@@
 Our Reducio-VAE is trained on a 400K video dataset collected from Pexels***https://www.pexels.com/.
-Pexels contains a large number of free and high-quality stock videos, where each video is accompanied with a short text description.
+Pexels contains a large number of free and high-quality stock videos, where each video is accompanied by a short text description.
 We train Reducio-VAE from scratch on $256\times 256$ videos with 16 FPS.
@@
 Specifically, our model significantly outperforms state-of-the-art 2D VAE, e.g., SD2.1-VAE and SDXL-VAE, by more than 5 db in PSNR.
-Moreover, Reducio-VAE also demonstrates better performance compared with VAEs designed for video in recent literature, e.g., OmniTokenizer and OpenSora-1.2.
+Moreover, Reducio-VAE also performs better than VAEs designed for video in recent literature, e.g., OmniTokenizer and OpenSora-1.2.
 Our VAE also outperforms the concurrent work, Cosmos-VAE, which also adopts a further compressed video latent space, by 0.2 in SSIM and 5 db in PSNR with an 8×8$\times$8 × higher down-sampling factor.
 We observe Reducio-VAE has worse rFVD than others on UCF-101.
-By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence result in a lower recontsruction quality.
-In contrast, Reducio-VAE places first on rFVD on pexels, which consists videos with much improved viusal quality and much higher resolution.
+By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence result in a lower reconstruction quality.
+In contrast, Reducio-VAE places first on rFVD on pexels, which consists of videos with much improved visual quality and much higher resolution.
 Reducio-DiT achieves a promising balance between efficiency and performance.
@@
 While DynamicCrafter and SVD-XT both fail to generate stable frames consistent with the given content frame, Reducio-DiT yields high visual quality with facial details in the content frame uncorrupted.
-Moreover, Reducio-DiT significantly boost the throughput of video LDMs.
+Moreover, Reducio-DiT significantly boosts the throughput of video LDMs.
 The compressed latent space facilitates Reducio-DiT to output $256^{2}$ videos of 16 frames in almost 1 second, which realize a $\mathbf{5.3\times}$ speed up compared with Lavie.
 Moreover, 16-frame $1024^{2}$ video generation costs Reducio-DiT only for 15.4 seconds on average, leading to a striking speedup of $\mathbf{16.6\times}$ over Lavie.
-We argue the core philosophy of Reducio lies in the extremely compressed latent space, which acts as the key role in accelerating both training and inference speed of video LDMs.
+We argue the core philosophy of Reducio lies in the extremely compressed latent space, which acts as the key role in accelerating both the training and inference speed of video LDMs.
 4.3 Ablation Studies
@@
 5.
-As for add-based fusion, we feed the content frame features into a convolution and replicate the output on temporal dimension.
+As for add-based fusion, we feed the content frame features into a convolution and replicate the output on the temporal dimension.
 Then we add video features with the content feature, together with a temporal embedding.
 Concerning linear-based fusion, we view the content condition as a single-frame video feature and concatenate it with the video feature by the temporal dimension.
-The concatenated features are then refined by a 3D convolution layer.
+A 3D convolution layer then refines the concatenated features.
 As observed in Tab.
@@
 We believe in the potential of adapting our works for longer video generation and will explore it as the future work.
+Reducio is purely a research project.
+Currently, we have no plans to incorporate Reducio into a product or expand access to the public.
+We will also put Microsoft AI principles into practice when further developing the models.
+In our research paper, we account for the ethical concerns associated with text-image-to-video research.
+To mitigate issues associated with training data, we have implemented a rigorous filtering process to purge our training data of inappropriate content, such as explicit imagery and offensive language, to minimize the likelihood of generating inappropriate content.
+Appendix A Implementation details
+Reducio-VAE.
+We demonstrate the detailed architecture of Reducio in Fig.
+8.
+Generally, we follow the training strategies of SD-VAE [40].
+We employ a tailored version of PatchGAN [18] based on 3D convolutions and optimize the model with $L_{1}$ loss, KL loss, perceptual loss [68], and GAN loss.
+While initializing the 2D encoder and the 3D VAE with SD-VAE pre-trained weights accelerates convergence, we find that freezing the 2D encoder leads to worse performance than training the full parameters.
+We follow LaMD [16] to feed the motion latent into a normalization layer to obtain the output of the VAE encoder.
+Therefore, in the stage of diffusion training, we use a scale factor of 1.0, which is multiplied by the input latent as inputs into DiT.
+Figure 8: The detailed architecture of Reducio-VAE with $f_{t}=4,f_{s}=32$.
+Table 9: Hyperparameters for Reducio-VAE
+Reducio-VAE
+$z$-shapeon $256^{2}$ videos $4\times 8\times 8\times 16$ $8\times 8\times 8\times 16$
+Channels (3d) 128
+Channels (2d) 128
+Ch Multiplier (3d) 1,2,2,4,4,4
+Ch Multiplier (2d) 1,2,2,4,4
+Depth 2
+Batch size 32 24
+Learning rate 4e-5
+Iterations 1,000,000
+Table 10: Comparison with more SOTA models on Vbench.
+Model Quality Semantic Total
+Score Score Score
+Show-1 [66] 80.42 72.98 78.93
+Lavie [53] 78.78 70.31 77.08
+VideoCrafter [5] 81.59 72.22 79.72
+OpenSora v1.2 [30] 81.35 73.39 79.76
+Lavie-2 [53] 83.24 75.76 81.75
+Pyramid Flow [19] 84.74 69.62 81.72
+VideoCrafter-2 [12] 83.27 76.73 81.97
+Reducio-DiT 82.24 78.00 81.39
+During inference, We split videos with a resolution over $256\times 256$ into overlapping spatial tiles, we fuse the encoded latent as well as the video output, in a similar manner with Movie Gen [35].
+Note that since Reducio-VAE employs a spatial down-sampling factor of 32, Reducio-VAE can only process video inputs whose width and height are divisible by 32.
+Otherwise, videos should be padded to meet this requirement before being fed into the VAE.
+Reducio-DiT.
+We elaborate on the details of content image conditions in Fig.
+9.
+During inference, we use classifier-free guidance [13] for better generation quality and set the default scale to 2.5.
+During training, we randomly drop image conditions at a probability of 0.1, as well as drop 10% text conditions.
+Figure 9: The overview of the efficient content condition solution for Reducio-DiT on high-resolution videos.
+Table 11: Detailed quantitative comparison with SOTA for text-to-video generation models on VBench.
+Model subject background temporal motion dynamic aesthetic imaging object multiple human color spatial scene appearance temporal overall
+consistency consistency flickering smoothness degree quality quality class objects action relationship style style consistency
+Lavie [53] 91.41 97.47 98.30 96.38 49.72 54.94 61.90 91.82 33.32 96.80 86.39 34.09 52.69 23.56 25.93 26.41
+Show-1 [66] 95.53 98.02 99.12 98.24 44.44 57.35 58.66 93.07 45.47 95.60 86.35 53.50 47.03 23.06 25.28 27.46
+VideoCrafter [5] 95.10 98.04 98.93 95.67 55.00 62.67 65.46 78.18 45.66 91.60 93.32 58.86 43.75 24.41 25.54 26.76
+OpenSora v1.2 [30] 96.75 97.61 99.53 98.50 42.39 56.85 63.34 82.22 51.83 91.20 90.08 68.56 42.44 23.95 24.54 26.85
+Lavie-2 [53] 97.90 98.45 98.76 98.42 31.11 67.62 70.39 97.52 64.88 96.40 91.65 38.68 49.59 25.09 25.24 27.39
+Pyramid Flow [19] 96.95 98.06 99.49 99.12 64.63 63.26 65.01 86.67 50.71 85.60 82.87 59.53 43.20 20.91 23.09 26.23
+VideoCrafter-2 [12] 97.17 98.54 98.46 97.75 42.50 65.89 70.45 93.39 49.83 95.00 94.41 64.88 51.82 24.32 25.17 27.57
+Reducio-DiT 98.05 99.13 98.45 98.77 27.78 64.02 67.67 91.49 69.91 92.60 89.06 52.85 54.90 25.16 26.40 28.87
+Appendix B Additional results
+We display the detailed performance comparison on VBench [17] in Tab.
+11 and Tab.
+10.
+As the results show, Reducio-DiT achieves a promising semantic score of 78.00, beating a range of state-of-the-art LDMs.